Please solve for X and please show all work.
15,000 = (400,000 + X)/40

Answer by stanbon(75887)

15,000 = (400,000 + X)/40
------
400,000 + x = 40*15000
----
x = 40*15000 - 400,000
----
x = 200,000
=================
Cheers,
Stan H.